OpenStudy (anonymous):

write as a radical expression 27 ^-2/3

OpenStudy (anonymous):

\[\sqrt[3]{27}^{-2}\]

OpenStudy (saifoo.khan):

\[\frac{1}{27^\frac23}\to \frac{1}{9}\]

OpenStudy (alexwee123):